3. One of the top-selling items in the container group at the museum’s gift shop is a birdfeeder. Sales are 18 units per week, and the supplier charges $60 per unit. The cost of placing an order with the supplier is $45. Annual holding cost for each unit is 25 percent of a feeder’s value, and the museum operates 52 weeks per year. Management currently chose to order 390 units when they place an order. What is the annual cost of the current ordering policy? How much can the shop save in terms of the annual cost if they order optimally? Assume that there is no lead time and no uncertainty in the sales volume.
Answer #1

EOQ= ((2*Annual Demand*Cost per order)/Holding cost)1/2

Demand= 18 pairs per week = 18*52 = 936 units annually

Order cost= 45$

Holding cost= .25*60= 15$

EOQ= (2*936*45/15)1/2 = 74.94 = 75 units

Total cost = Ordering cost + holding cost

= (Annual demand/EOQ)*ordering cost + (EOQ/2)*Cost per pair*.40

Current quantity = 390 units

= (936/390)*45 + (390/2)*15 = 108 + 2925

= 3033 $

EOQ = 75 units

Total cost = = (936/75)*45 + (75/2)*15 = 561.6 + 562.5

= 1124.1 $

Saving with EOQ quantity = 3033 - 1124.1

= 1908.9 $